OpenStudy (anonymous):

simplify the inverse trig expression with answers [0,2pi) (more in comments)

OpenStudy (anonymous):

\[\sin^{-1} (-1/2)\]

OpenStudy (anonymous):

how many answers do you think you'll get here?

OpenStudy (anonymous):

sin(what angle) = -1/2.... the y-coordinate is -1/2..

OpenStudy (anonymous):

@kaisan , look at your unit circle.... what point has it's y-coordinate -1/2?

OpenStudy (anonymous):

here.. page 3 http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

because the y coordinate is negative the angle must be in the 3rd and 4th quadrants.

OpenStudy (anonymous):

yup i see it ^^ sorry for the long reply

OpenStudy (anonymous):

i just get confused when i see \[\sin^{-1} \]

OpenStudy (anonymous):

yeah, when i see that it change it to this question: sin (what angle) = -1/2 that's all it means...

OpenStudy (anonymous):

ok ^^ thx again